OpenStudy (anonymous):

What is the surface area of a cube with side length 9 km?

OpenStudy (anonymous):

a=9 \[surfaceareaofacube=6a^2\]

OpenStudy (anonymous):

total surface area?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ok just plug in what mathpro wrote and u will get ur answer: 6(9^2)=

OpenStudy (anonymous):

so whats the answer cause i got 108 and its not a chocie

OpenStudy (anonymous):

but your answer is wrong try once more!!! 9^2=9*9=?

OpenStudy (anonymous):

can u just give me the answerr!

OpenStudy (anonymous):

9^2=81 81*6=486

OpenStudy (anonymous):

u did 6*2*9 u need to do 6*9*9